how many pounds of nuts selling for 6 dollars per lb and raisins selling for 3 dollars per lb should a person combined to obtain 180 lb of a trail mix selling for 4 dollars per lb
Answer by stanbon(75887)

how many pounds of nuts selling for 6 dollars per lb and raisins selling for 3 dollars per lb should a person combined to obtain 180 lb of a trail mix selling for 4 dollars per lb
------------------
Equations:
Quantity Eq:: n + r = 180 lb
Value Eq:: 6n + 3r = 4*180
-----
n + r = 180
2n+ r = 4*60
-------
Subtract and solve for "n":
n = 240-180
n = 60 lb (lbs of nuts needed)
------------------
r = 180-60
r = 120 lbs (lbs of raisins needed)
------------------
Cheers,
Stan H.
-------------